# Don’t Fine-Tune, Decode: Syntax Error-Free Tool Use via Constrained Decoding

### v2

@@ 1 Introduction
We evaluate ToolDec by applying it to 5 base models and evaluating it on 4 benchmarks. On base models that are specifically fine-tuned for tool use purposes, such as ToolLLM Qin et al. (2023) and ToolkenGPT Hao et al. (2023), ToolDec significantly improves their accuracy by at most 21 points and greatly reduce the number of tokens in the output. On base models that initially have zero accuracy on ToolBench Qin et al. (2023) such as Mistral Jiang et al. (2023) and Vicuna Chiang et al. (2023), ToolDec improves their accuracy from 0 to at most 60%.
@@ 2 Related Work
Constrained decoding and finite-state machines. Previous constrained decoding methods reduce the large search space of lexically constrained decoding with finite-state machines (Anderson et al., 2017), grouping together similar candidates (Hokamp & Liu, 2017), and better search algorithms (Miao et al., 2019; Lu et al., 2021, 2022). However, lexical constraints are not expressive enough to regulate tool calls. While finite-state machines have to be weighted and probabilistic to deal with the soft constraints in natural language (Eisner, 2002; Rastogi et al., 2016), the constraints for syntactic tool calls are hard constraints that are much easier for FSMs. Grammar-constrained decoding has been used for structural NLP tasks such as code generation Yin & Neubig (2017), semantic parsing Stengel-Eskin et al. (2024), coreference resolution, POS tagging and many others Geng et al. (2023). Using external tools is different from these tasks because the syntax and semantics of tools are unknown to the model until inference time.

### v3

@@ 1 Introduction
We evaluate ToolDec by applying it to 5 base models and evaluating it on 4 benchmarks. On base models that are specifically fine-tuned for tool use purposes, such as ToolLLM (Qin et al., 2023) and ToolkenGPT (Hao et al., 2023), ToolDec significantly improves their accuracy by at most 21 points and greatly reduce the number of tokens in the output. On base models that initially have zero accuracy on ToolBench (Qin et al., 2023) such as Mistral (Jiang et al., 2023) and Vicuna (Chiang et al., 2023), ToolDec improves their accuracy from 0 to at most 60%.
@@ 2 Related Work
Constrained decoding and finite-state machines. Previous constrained decoding methods reduce the large search space of lexically constrained decoding with finite-state machines (Anderson et al., 2017), grouping together similar candidates (Hokamp & Liu, 2017), and better search algorithms (Miao et al., 2019; Lu et al., 2021, 2022). However, lexical constraints are not expressive enough to regulate tool calls. While finite-state machines have to be weighted and probabilistic to deal with the soft constraints in natural language (Eisner, 2002; Rastogi et al., 2016), the constraints for syntactic tool calls are hard constraints that are much easier for FSMs. Grammar-constrained decoding has been used for structural NLP tasks such as code generation (Yin & Neubig, 2017), semantic parsing (Stengel-Eskin et al., 2024), coreference resolution, POS tagging and many others (Geng et al., 2023). Willard & Louf (2023) uses finite state machines to guide LLMs to generate outputs efficiently and conform to grammar.